A Foucault pendulum is used to demonstrate the rotation of the Earth. They can usually be found in science museums. A certain pendulum has a 120-kg bob swinging on a 16.2-m-long wire with a maximum amplitude of 6.3°. Show all work.

A. Sketch the pendulum at its maximum displacement and label all known quantities.

B. What is the difference in gravitational potential energy between the maximum displacement and the equilibrium point?

C. What is the maximum speed of the bob?

D. Where in the oscillation is the bob when it has its maximum speed?

E. What is the period and frequency of the oscillation?
Answer #1

b)

from geometry

cos x = L - h / L

16.2 cos 6.3 = 16.2 - h

h = 0.098 m

difference in gravitational potential energy

U = mgh = 120* 9.8* 0.098 = 11.256 J

=====

c)

as per energy conservation

0.5 m v^2 = mgh

0.5 v^2 = 9.8* 0.098

v = 1.386 m/s

=====

d)

max speed is at lowest point ( mean position)

=====

e)

T = 2pi sqrt ( L/g) = 2 pi sqrt ( 16.2 / 9.8)

T = 8.074 s

frequency,

f = 1/T = 0.1238 Hz

======

Comment before rate in case any doubt, will reply for sure.. goodluck